Use the following information to answer question 5 and 6 Suppose that the current spot exchange rate between Japanese Yen and Euro is ¥130/€ and the one-year forward exchange rate is ¥138.25/€. The one-year interest rate is 2.0 % in yens and 1.25% in euro.

5. According to the Interest Rate Parity condition, what is the 1 year forward exchange rate?

a. ¥139.27/€
b. ¥130.96/€
c. ¥129.04/€
d. ¥137.23/€

6. What is your arbitrage strategy if you can borrow 10 million Yen or 76,923.08 euro amount, at the current spot exchange rate?

a. €4,107.17
b. €73,779.39
c. ¥767,548.72
d. ¥567,548.72
Answer #1

As per IRPT, fair forward rate = spot rate*(1+Interest rate in Yen)/(1+Interest rate in Euro)

= 130(1+0.02)/(1+0.0125)

= Yen 130.96/Euro

i.e. b

4.Starting with yen 10,000,000

Convert into Euro at spot rate and get 10,000,000/130 = Euro 76,923.08

Invest and get 76,923.08(1+0.0125) = Euro 77,884.6185

Convert into Yen at forward rate and get 77,884.6185*138.25 = Yen 10,767,548.72

Pay back Loan 10,000,000(1+0.02) = Yen 10,200,000

Arbitrage Profit = Yen 567,548.72

i.e. d